Could someone please help me with this word problem, this problem needs to be solved by using systems of linear equations. Thank you in advance for your help.
In a triangle, the sum of the interior angles is 180 degrees. If angle A is three times as large as angle B, and angles B and C sum to 90 degrees, what is the measure of each angle?

Answer by stanbon(75887)

In a triangle, the sum of the interior angles is 180 degrees. If angle A is three times as large as angle B, and angles B and C sum to 90 degrees, what is the measure of each angle?
------------------
A + B + C = 180
A -3B + 0 = 0
0 + B + C = 90
------------------
Solve the system by any means to get:
A = 90 degrees
B = 30 degrees
C = 60 degrees
========================
Cheers,
Stan H.